# What Makes LLM Agent Simulations Useful for Policy? Insights From an Iterative Design Engagement in Emergency Preparedness

### v1

@@ Abstract.
There is growing interest in using Large Language Models as agents (LLM agents) for social simulations to inform policy, yet real-world adoption remains limited. This paper addresses the question: How can LLM agent simulations be made genuinely useful for policy? We report on a year-long iterative design engagement with a university emergency preparedness team. Across multiple design iterations, we iteratively developed a system of 13,000 LLM agents that simulate crowd movement and communication during a large-scale gathering under various emergency scenarios. These simulations informed actual policy implementation, shaping volunteer training, evacuation protocols, and infrastructure planning. Analyzing this process, we identify three design implications: start with verifiable scenarios and build trust gradually, use preliminary simulations to elicit tacit knowledge, and treat simulation and policy development as evolving together. These implications highlight actionable pathways to making LLM agent simulations that are genuinely useful for policy.
@@
Figure 1. We report on an iterative design engagement with emergency preparedness policymakers to develop LLM agent simulations that inform policy implementation. Through cycles of design and validation, simulations shifted from being distrusted to trusted, and were ultimately adopted in policy implementation. This usefulness did not follow a linear path from technical sophistication to institutional adaptation, but instead emerged through an iterative, stakeholder-engaged design process. :
@@ 1. Introduction
Large language models (LLMs) have unlocked new possibilities for simulating human behavior (Park et al., 2023, 2024; Binz et al., 2025; Xie et al., 2024). Beyond generating text, LLMs can be deployed as agents, capable of reasoning, conversing, and interacting with defined environments (Sumers et al., 2023; Li et al., 2023; Du et al., 2024). When deployed at scale, collections of such agents have been shown to reproduce emergent social phenomena, from collective decision-making to information propagation (Piao et al., 2025; Gao et al., 2023; Cui et al., 2025). These findings have inspired a growing thread of research on LLM agent simulation systems, promoted as tools to explore the potential impacts of policies in silico before they unfold in practice (Li et al., 2024a; Horton, 2023; Karten et al., 2025; Zhang et al., 2024; Jin et al., 2025; Steenstra et al., 2025; Williams et al., 2023; Li et al., 2024b; Liu et al., 2025b, a). Proponents argue that such systems could enable stakeholders to experiment with alternative scenarios and anticipate possible consequences (Hwang et al., 2025; Anthis et al., 2025; Xu et al., 2023; Ma et al., 2024).
Yet despite this promise, no policymaking body has, to our knowledge, adopted these systems in decision-making. Instead, they remain largely confined to academic demonstrations, with limited real-world validation and little integration into the workflows, priorities, and institutional constraints of policymakers. Critics caution that LLM agents can risk reproducing stereotypes, amplifying biases, or offering “black-box” outputs that policymakers should not trust (Arnstein, 1969; Agnew et al., 2024; Li et al., 2025), while others warn that they may foster false confidence by encouraging decision-makers to treat speculative outputs as predictive truths (Kapania et al., 2025; Zhou et al., 2024).
These gaps highlight a central tension: while LLM agent simulations hold potential as exploratory tools, their current form falls short of the institutional legitimacy and trustworthiness required for real-world policy use. The tradition of Human-Computer Interaction (HCI) suggests a possible pathway forward, with its rich history of participatory and user-centered design methodologies. These approaches emphasize the involvement of stakeholders in technology development to ensure systems align with actual practices and needs, rather than abstract potential (Schuler and Namioka, 1993; Abras et al., 2004; Sanders and Stappers, 2008).
Building on this tradition, we ask: How can LLM agent simulations be made genuinely useful for policy? The “last-mile” usefulness of these simulations can hinge less on advances in models’ capabilities or frameworks’ sophistication and more on the design process itself: engaging policymakers iteratively, grounding simulations in their lived challenges, validating outputs against real-world evidence, and collaboratively exploring how simulation results can inform concrete implementation proposals.
@@
In this paper, we report on a year-long iterative design engagement with a university emergency preparedness and response team—the institutional decision-makers responsible for preparedness policy implementation (hereafter referred to as the “policymakers”). We began with semi-structured interviews to map their existing workflows and decision-making practices, and to probe how simulation might fit within them. We eventually identified graduation commencement—a large, complex gathering event involving thousands of students, families, and staff—as a concrete policy domain in which evacuation preparedness was both pressing and open to innovation. Over five iterative design phases, we developed and refined an LLM agent simulation system to generate crowd movement and communication dynamics under varying conditions of scale, physical constraints, and information scarcity. At each step, we incorporated policymakers’ feedback, validated the simulator against ground-truth data, and worked toward actionable proposals for policy implementation refinement.
@@
Alongside these design iterations, we analyzed our design process to surface broader insights into when and why simulations were perceived as relevant, trustworthy, and actionable by policymakers. Our analysis suggests three key design implications for making LLM agent simulations useful for policy implementation. First, start with verifiable scenarios and build trust gradually. Policymakers only engage seriously with simulations when at least part of the output can be checked against reality. Beginning with routine, recognizable cases helps establish credibility, which can then support exploration of more speculative situations. Second, use preliminary simulations to elicit tacit knowledge. Even imperfect or “wrong” outputs can be productive: they prompt policymakers to surface the unstated practices and contextual details—such as roles, relationships, or environmental factors—that matter most for policy implementation. Third, treat simulation and policy development as evolving together. Simulation capabilities and institutional requirements rarely align from the start. As simulations mature, they shape policymaker expectations, and as policymaker expectations shift, they reshape simulation needs. Framing both as evolving together enables simulations to become durable tools that support decision-making in dynamic institutional contexts shaped by emerging AI capabilities.
@@
    Iterative design of LLM agent simulations in emergency preparedness: We show how collaborative engagement enabled simulations to evolve from academic prototypes into tools that informed concrete preparedness practices.
@@
    Insights on simulation-to-policy relevance: We distill five insights that determine why and how LLM agent simulations are perceived as useful and adopted by policymakers.
@@
    Design implications for building useful LLM agent simulations for policy implementation: We provide concrete guidance that developers and designers can apply when developing simulations to inform policy implementation.
Together, these contributions demonstrate not only the potential of LLM agent simulations but also the conditions under which they can move beyond academic demonstration toward institutional uptake.
@@ 2.1. LLM Agent Simulations and Social Modeling
LLM agents have been shown to produce strikingly human-like behaviors when situated in simulated environments. Early demonstrations revealed that agents equipped with memory and reflection can spontaneously form coalitions, spread invitations, and coordinate collective events (Park et al., 2023, 2024). Building on this foundation, researchers have extended the paradigm to diverse domains: clinical workflows in hospitals, classroom dynamics in education, macroeconomic trading, and other large-scale social processes (Binz et al., 2025; Xie et al., 2024; Sumers et al., 2023; Li et al., 2023; Du et al., 2024). Multi-agent systems have further been used to explore policy-relevant questions, from epidemic response to online misinformation and environmental decision-making (Piao et al., 2025; Gao et al., 2023; Cui et al., 2025; Williams et al., 2023; Li et al., 2024a; Liu et al., 2025b; Zhang et al., 2024; Steenstra et al., 2025; Liu et al., 2025a; Hwang et al., 2025; Anthis et al., 2025; Xu et al., 2023; Ma et al., 2024). Across these cases, the appeal lies in the promise of “in silico” policy testing—running counterfactual scenarios at scale that would be costly or unethical to attempt in the real world (Horton, 2023; Karten et al., 2025; Li et al., 2024b; Jin et al., 2025).
@@
2.2. HCI and Policy Implementation
HCI has long been concerned with civic technologies and decision-support systems that shape governance and public policy. Early work introduced platforms for participatory urban planning and neighborhood decision-making (Healey et al., 2008; Abdalla et al., 2016). More recent systems like PolicyCraft structured deliberations to improve consensus and justification among stakeholders (Kuo et al., 2025). Studies in crisis informatics highlight similar goals in emergency response, where tools support sensemaking and communication among responders and the public (Palen and Liu, 2007; Vieweg et al., 2010; Starbird and Palen, 2011).
@@ 2.3. Engaging Stakeholders in Design
To navigate such complexities, HCI researchers have developed a wide repertoire of stakeholder engagement methods. Traditions of participatory design, user-centered design, and community-based co-design have demonstrated how iterative cycles of prototyping, feedback, and negotiation can surface hidden needs and reconcile value tensions (Schuler and Namioka, 1993; Kuo et al., 2025; Abras et al., 2004; Sanders and Stappers, 2008; Han et al., 2024; Chen et al., 2025; Holstein et al., 2019a; Kawakami et al., 2024; Hutchinson et al., 2003b; Weitz et al., 2024; Hutchinson et al., 2003a). Longitudinal collaborations with municipal offices (Corbett and Le Dantec, 2018), community organizations (Calvo and De Rosa, 2017), and health institutions (Slattery et al., 2020; O’Brien et al., 2021) underscore the importance of partnerships that adapt alongside institutional contexts. Agile and iterative approaches have also been adopted in high-stakes domains such as education, health, and emergency response to foster responsiveness and resilience (Shore and Warden, 2021; Abrahamsson et al., 2017; Thomer et al., 2016).
With the rise of AI-powered systems, however, new challenges emerge. Unlike traditional tools, AI outputs are probabilistic, opaque, and sometimes unpredictable, complicating stakeholder understanding and co-design (Abdul et al., 2018; Dolata et al., 2024). Recent work has explored participatory approaches to algorithm design to surface values, fairness concerns, and accountability requirements (Deng et al., 2025; Holstein et al., 2019b), while others emphasize new representational tools to make AI legible in design workshops (Lindley et al., 2020; Ma’ayan et al., 2020). Together, these efforts point toward a growing consensus: longitudinal, participatory engagements are critical not only for uncovering requirements, but also for fostering trust, negotiating unpredictability, and embedding AI systems into institutional practice. This study contributes to these engagement traditions by addressing the specific challenges of LLM agent simulations for emergency preparedness policy implementation, where outputs are both generative and collective, and where contextual requirements are often tacit.
@@ 3. Policy Domain: Emergency Preparedness
To explore what makes LLM agent simulations useful for policy, we identified emergency preparedness as our design context. Emergency preparedness represents a uniquely fitting domain because real-world emergency experiments are practically and ethically impossible, observational data from crises are scarce and difficult to generalize, and policymakers must make high-stakes decisions under profound uncertainty (Helbing et al., 2005, 2006). Unlike domains where interventions can be validated through controlled studies, emergency preparedness requires anticipating rare, unpredictable events with limited empirical evidence. LLM agent simulations might offer a way to address these challenges by enabling exploration of evacuation dynamics, testing communication protocols, and examining crowd behaviors under conditions that cannot be safely replicated in reality.
@@
Building on prior work and existing practices, we hypothesized that—given the potential and limitations of state-of-the-art LLMs—the usefulness of LLM agent simulations in policy contexts hinges less on advances in modeling fidelity or computational scale, and more on the design process. Simulations can become valuable when they are developed through iterative engagement with policymakers, grounded in their lived challenges, validated against real-world evidence, and collaboratively explored as tools for shaping concrete policy proposals.
@@
To situate this inquiry, we partnered with a university’s emergency preparedness and response team. This team is responsible for developing and implementing emergency preparedness policies across diverse scenarios—from preparing evacuation procedures for large-scale gatherings such as commencement and orientation, to coordinating responses to natural disasters (e.g., tornadoes, floods), and managing high-stakes emergencies such as active shooter incidents. Because emergencies are inherently rare and information-poor, the team must act under profound uncertainly—balancing institutional policy, safety concerns, and communication clarity while coordinating across stakeholders such as university administration, student groups, local police, and nearby universities.
@@
A central tool in this process is the after-action report. After each major university-wide event, the team documents both their preparations beforehand and their actions during the events. These reports highlight improvements over prior processes, identify gaps, and record strategies that worked in practice. The team reviews past reports to refine future policy implementation, the university sometimes adopts successful practices from these reports. In this way, after-action reporting functions as a bridge between day-of response and long-term institutional change, making it a critical site for understanding how simulations might inform policy implementation.
@@
We began our partnership by introducing prior work on model-based emergency coordination and exploring how it could be extended with LLMs. At that stage, the policymaker team already understood both the promise and limitations of social simulations for their domain. They also recognized the natural-language generation capabilities of LLMs as potentially useful for addressing pressing challenges in emergency communication and social media management. Several members expressed particular interest in generating realistic variations of agent behavior through persona-based prompting techniques, seeing this as a way to capture community diversity and stress-test communication strategies. At the same time, the team made no commitment to incorporate simulation outcomes into their policy implementation, underscoring that the collaboration was exploratory rather than prescriptive.

### v2

# What Makes LLM Agent Simulations Useful for Policy Practice? An Iterative Design Study in Emergency Preparedness

@@ Abstract.
Policymakers must often act under conditions of deep uncertainty, such as emergency response, where predicting the specific impacts of a policy apriori is implausible. Large Language Model (LLM) agent simulations have been proposed as tools to support policymakers under these conditions, yet little is known about how such simulations become useful for real-world policy practice. To address this gap, we conducted a year-long, stakeholder-engaged design process with a university emergency preparedness team. Through iterative design cycles, we developed and refined an LLM agent simulation of a large-scale campus gathering, ultimately scaling to 13,000 agents that modeled crowd movement and communication under various emergency scenarios. Rather than producing predictive forecasts, these simulations supported policy practice by shaping volunteer training, evacuation procedures, and infrastructure planning. Analyzing these findings, we identify three design process implications for making LLM agent simulations that are useful for policy practice: start from verifiable scenarios to bootstrap trust, use preliminary simulations to elicit tacit domain knowledge, and treat simulation capabilities and policy implementation as co-evolving.
@@
Figure 1. An iterative design engagement with emergency preparedness policymakers developed LLM agent simulations to inform policy implementation. Through cycles of design and validation, simulations shifted from being distrusted to trusted and were ultimately integrated into preparedness practice. This usefulness did not follow a linear path from technical sophistication to institutional adaptation, but instead emerged through an iterative, stakeholder-engaged design process.
Diagram showing how developers and policymakers iteratively co-design large language model agent simulations for emergency preparedness. Developers create simulations that policymakers review, moving perceptions from useless to useful and from distrust to trust. A stadium evacuation simulation illustrates how policy changes can be tested, with a bar chart showing improved evacuation time under suggested processes.
@@ 1. Introduction
Helping policymakers reason and act under conditions of deep uncertainty—situations like emergency response, where outcomes depend on many unknowable environmental and social factors—is a longstanding challenge in decision-making research and human-computer interaction (Lempert, 2002; Waddell and Borning, 2004). Large language models (LLMs) offer a potential pathway for addressing this challenge: they have been used as interacting agents whose behavior is generated through natural language, enabling large-scale populations to exhibit emergent social behaviors—such as coordination, collective decision-making, and information propagation—grounded in interpretation and communication (Park et al., 2023; Sumers et al., 2023; Du et al., 2024; Piao et al., 2025; Gao et al., 2023). These capabilities have inspired a growing body of work on LLM agent simulation systems, which envision policymakers using these systems to experiment with alternative scenarios, communication strategies, or interventions before implementation (Horton, 2023; Karten et al., 2025; Hwang et al., 2025; Hou et al., 2025).
However, what it means for LLM agent simulations to be useful for policy remains unclear. Much existing work in such simulations implicitly frames policy in terms of prediction or optimization. In practice, policy work involves interpreting uncertain information, coordinating across roles, developing procedures, training personnel, and adapting resource allocations under institutional constraints—rather than selecting an optimal action based on forecasts alone (Lempert, 2002). At the same time, applying LLM agent simulations to policy contexts introduces new risks: they may reproduce stereotypes, amplify biases, or generate outputs that are easy to read but difficult to causally explain or defend within policy practice (Arnstein, 1969; Agnew et al., 2024; Li et al., 2025). These concerns foreground the need to understand the conditions under which LLM agent simulations can be meaningfully integrated into policy practice without causing harm.
This motivates a central question for this work: How can LLM agent simulations be made genuinely useful for policy practice? We address this question by treating policy not as a set of outcomes to be optimized, but as institutionally situated practice. We use the term “policy” to refer to both institutionally defined goals (e.g., protecting life and safety) and the concrete procedures, practices, and resource decisions through which those goals are enacted in everyday work. Accordingly, the challenge is to understand how LLM agent simulations can become useful within the constraints and routines of real-world policy implementation.
Rather than treating simulations as standalone analytical artifacts, the tradition of Human-Computer Interaction (HCI), grounded in participatory and user-centered design methodologies, suggests a design-oriented way forward. These approaches emphasize stakeholder-engaged design processes to ensure systems align with actual practices and needs, rather than on abstract technological potential (Schuler and Namioka, 1993; Abras et al., 2004; Sanders and Stappers, 2008). From this perspective, LLM agent simulations can be studied not only as prediction engines, but as interactive tools whose last-mile usefulness depends less on advances in model capability than on how they are iteratively developed with policymakers—grounded in lived challenges and collaboratively explored as part of concrete policy implementation work (Lempert, 2002).
@@
We report on a year-long iterative design engagement with a university emergency preparedness and response team—the institutional decision-makers responsible for preparedness policy implementation (hereafter referred to as the “policymakers”). We began with semi-structured interviews to understand existing workflows and decision-making practices, and points where simulation might productively intervene. Through this process, we identified graduation commencement—a large, complex gathering event involving thousands of students, families, and staff—as a concrete policy domain in which evacuation preparedness was both pressing and open to innovation. Traditionally, evacuation has primarily been modeled as a physics-based or geometric problem (Helbing et al., 2005). However, we found that preparedness policy depended critically on deeply social and institutional dimensions — e.g., communication, coordination, role expectations, and trust — that are not well captured by these simple physics-based models. Over five iterative design phases, we developed and refined an LLM agent simulation system to model crowd movement and communication dynamics under varying conditions of scale, physical constraints, and information scarcity. At each step, we incorporated policymakers’ feedback, validated the simulator against ground-truth data, and worked toward actionable proposals that could inform training, protocols, and the decision-making process itself.
@@
Alongside these design iterations, we analyzed the design process to understand when and why simulations were perceived as relevant, trustworthy, and actionable by policymakers. From this analysis, we distill three design process implications for making LLM agent simulations useful for policy practice. First, start with verifiable scenarios to build trust gradually; policymakers engage meaningfully only when their focal aspects of simulation outputs can be checked against reality. Second, preliminary simulations — even when imperfect — can serve as productive technology probes (Hutchinson et al., 2003) that elicit tacit knowledge about roles, relationships, and contextual factors that matter in policy implementation. Third, simulation and policy development should be treated as interdependent and evolving together; as simulations mature, they reshape policymaker expectations, and as institutional priorities shift, they redefine simulation requirements.
@@
    We present an in-depth account of the iterative design of an LLM agent simulation system for emergency preparedness, demonstrating how sustained collaboration enabled the system to evolve from academic prototypes into a tool that informed concrete preparedness practices  [^1] The core contribution of this work is not the system itself, but the design insights we uncovered in using the system as a technology probe..
@@
    We provide insights into simulation-to-policy relevance by articulating why and how LLM agent simulations come to be perceived as useful, trustworthy, and actionable within institutional policy contexts.
@@
    We offer design process implications that provide actionable guidance for researchers and practitioners seeking to develop LLM agent simulations that support policy practice through stakeholder-engaged design.
Together, these contributions articulate the conditions under which LLM agent simulations can move beyond mere predictive calculators to iterative, stakeholder-engaged tools, and become meaningfully integrated into policy practice.
@@ 2.1. LLM Agent Simulations and Social Modeling
Simulating human behavior in complex environments has traditionally relied on ABMs (Schelling, 1971; An, 2012; Macal and North, 2005). They have proven effective for modeling physical dynamics, such as crowd flow, geometric bottlenecks, and collision avoidance in evacuation scenarios (Joo et al., 2013; Rozo et al., 2019; Chen and Zhan, 2008; Helbing and Molnar, 1995). However, these models often rely on fixed, pre-defined rules that struggle to capture the ambiguity and relational complexity of human decision-making during crises (Sun et al., 2016; DeAngelis and Diaz, 2019; Kostakos et al., 2010; Aguirre et al., 2011). Real-world emergencies are rarely reducible to physics problems; they are social events dominated by information scarcity, conflicting social roles, and the need to interpret vague communication. In addition, accuracies of ABMs often highly depend on the parameters set by their users, making it hard for novices to flexibly tune for new scenarios (Groeneveld et al., 2017; Ligmann-Zielinska et al., 2014; Cheng et al., 2018).
LLM agents offer a promising alternative by introducing semantic reasoning and generative behavior into social simulation. Because LLM agents are generative and draw on broad knowledge, they have the potential to address shortcomings of traditional ABMs, including limited ability to model complex human dynamics and heavy dependence on manually tuned parameters (Gao et al., 2024; Gurcan, 2024). Collection of these agents in social contexts have been shown to spontaneously form coalitions, spread invitations, and coordinate collective events (Park et al., 2023, 2024). Researchers have extended this paradigm to diverse domains, including clinical workflows, classroom dynamics, and macroeconomic trading (Binz et al., 2025; Xie et al., 2024; Sumers et al., 2023; Li et al., 2023; Du et al., 2024). In policy-relevant domains, multi-agent systems have been used to explore epidemic response, misinformation spread, and environmental decision-making (Hwang et al., 2025; Li et al., 2024a; Xu et al., 2023; Ma et al., 2024; Hou et al., 2025). The appeal of these systems lies in “in silico” policy testing—running counterfactual scenarios at scale that would be costly or unethical to attempt in the real world (Horton, 2023; Karten et al., 2025; Li et al., 2024b; Jin et al., 2025).
@@
2.2. HCI and Public Policy
HCI has long been concerned with civic technologies and decision-support systems that shape governance and public policy. Here in this paper, public policy refers to the processes of implementation through which stable institutional goals are translated into organizational routines, tools, and decisions, rather than to formal rules or legislation alone. Early work introduced platforms for participatory urban planning and neighborhood decision-making (Healey et al., 2008; Abdalla et al., 2016). More recent systems like PolicyCraft structured deliberations to improve consensus and justification among stakeholders (Kuo et al., 2025). Studies in crisis informatics highlight similar goals in emergency response, where tools support sensemaking and communication among responders and the public (Palen and Liu, 2007; Vieweg et al., 2010; Starbird and Palen, 2011).
@@ 2.3. Engaging Stakeholders in Design
To navigate such complexities, HCI researchers have developed a wide repertoire of stakeholder engagement methods. Traditions of participatory design, user-centered design, and community-based co-design have demonstrated how iterative cycles of prototyping, feedback, and negotiation can surface hidden needs and reconcile value tensions (Schuler and Namioka, 1993; Abras et al., 2004; Sanders and Stappers, 2008; Hutchinson et al., 2003). Longitudinal collaborations with municipal offices (Corbett and Le Dantec, 2018), community organizations (Calvo and De Rosa, 2017), and health institutions (Slattery et al., 2020; O’Brien et al., 2021) underscore the importance of partnerships that adapt alongside institutional contexts. Agile and iterative approaches have also been adopted in high-stakes domains such as education, health, and emergency response to foster responsiveness and resilience (Shore and Warden, 2021; Abrahamsson et al., 2017; Thomer et al., 2016).
With the rise of AI-powered systems, however, new challenges emerge. Unlike traditional tools, AI outputs are probabilistic, opaque, and sometimes unpredictable, complicating stakeholder understanding and co-design (Abdul et al., 2018; Dolata et al., 2024). Recent work has explored participatory approaches to algorithm design to surface values, fairness concerns, and accountability requirements (Deng et al., 2025; Holstein et al., 2019), while others emphasize new representational tools to make AI legible in design workshops (Lindley et al., 2020; Ma’ayan et al., 2020). Together, these efforts point toward a growing consensus: longitudinal, participatory engagements are critical not only for uncovering requirements, but also for fostering trust, negotiating unpredictability, and embedding AI systems into institutional practice. This study contributes to these engagement traditions by addressing the specific challenges of LLM agent simulations for emergency preparedness policy implementation, where outputs are both generative and collective, and where contextual requirements are often tacit.
@@ 3. Policy Domain: Emergency Preparedness
To explore what makes LLM agent simulations useful for policy practice, we identified emergency preparedness as our design context (Palen et al., 2010). Emergency preparedness is a particularly fitting domain because field experiments are practically and ethically impossible, observational data are scarce and difficult to generalize, and policymakers must make high-stakes decisions under uncertainty — where the timing, form, and consequences of emergencies cannot be reliably predicted in advance (Helbing et al., 2005, 2006). Unlike domains where interventions can be validated through controlled studies, emergency preparedness requires anticipating rare, unpredictable events with limited empirical grounding (Comfort, 2007). In practice, policy in emergency preparedness is often enacted through concrete, operational decisions rather than strategic or legislative choices. Policymakers must decide how many coordinators or volunteers to deploy, where to position them, which exits to prioritize under different threat scenarios, how to configure accessible seating, and how to frame announcements that guide crowd movement during evacuation. These decisions are highly interdependent, context-specific, and they are difficult to evaluate in advance. Yet they fundamentally shape how emergency plans are implemented on the ground. Under such conditions, simulations are less effective when treated as traditional policy analysis tools that prioritize ranking policy actions based on predicted outcomes (Morgan et al., 1990). Instead, they can be valued for their ability to support exploration, reflection, and deliberation under uncertainty—helping policymakers examine assumptions, surface blind spots, and reason through possible responses before crises occur (Lempert, 2002).
In addition, many challenges in contemporary emergency preparedness are increasingly shaped by collective communication dynamics rather than individual physical movement alone (Castillo, 2016; Reuter and Kaufhold, 2018). While early research emphasized physics-based movement, it is now understood that physical and social factors interact, with group processes on communication shaping evacuation performance (Nilsson and Johansson, 2009; van der Wal et al., 2021; Templeton et al., 2024). As digital and online technologies have become central to how people receive, interpret, and respond to emergency information (Starbird et al., 2014; Takayasu et al., 2015; Templeton et al., 2023), policy effectiveness depends not only on evacuation routes or crowd flows, but also on how messages are framed, how trust is established, and how individuals coordinate in response to uncertain and evolving information (Thompson et al., 2017; Shirado et al., 2020). These dynamics are inherently social and semantic, involving language, interpretation, emotion, and role expectations that unfold through interaction (Drabek, 2012; Utz et al., 2013; Jia et al., 2021). This shift complicates emergency preparedness for conventional policy practices and reduces the utility of tools that assume static or physics-based human behavior. At the same time, it makes emergency preparedness a particularly relevant context for examining LLM-based agent simulations, which can represent communicative behavior and meaning-making processes alongside physical action.
@@
Given this unpredictability and domain complexity, the usefulness of LLM agent simulations for emergency preparedness can hinge less on advances in modeling fidelity or computational scale, and more on the design process. In practice, simulations can become useful when they are iteratively developed with policymakers, grounded in their lived challenges, validated against real-world evidence, and collaboratively explored as tools for shaping concrete policy proposals (Moss, 2002).
@@
To situate this inquiry, we partnered with a university’s emergency preparedness and response team. This team is responsible for developing and implementing emergency preparedness policies across diverse scenarios, including large-scale gatherings such as commencement and orientation, natural disasters such as tornadoes and floods, and high-stakes emergencies such as active shooter incidents. Their policymaking practice was largely experience-driven and discussion-based, and did not involve the routine use of simulation tools. For recurring events, the team typically adopted plans from previous years as a baseline, using their collective professional judgment to identify incremental improvements. At the same time, because emergencies are inherently rare and information-poor, the team operated under substantial uncertainty, balancing institutional policy, safety concerns, and communication clarity while coordinating across stakeholders such as university administration, student groups, local police, and neighboring institutions.
@@
A central tool in this process is the after-action report. Following each major university-wide event, the team documents preparatory decisions and actions taken during the event, highlighting what worked, what failed, and where improvements are needed. These reports inform future policy implementation and occasionally lead to institution-wide changes. In this way, after-action reporting serves as a bridge between situated response and long-term organizational learning, making it a critical site for understanding how simulations might inform policy practice.
@@
We began our partnership by introducing prior work on model-based emergency coordination and exploring how it could be extended with recent advances in LLMs. At that stage, the policymaker team was already familiar with both the promise and limitations of simulations in their domain. For example, while they were aware of existing agent-based models for emergency preparedness, they had not incorporated such tools into their process, citing concerns about their ability to capture the situational detail and institutional realities relevant to their practice. At the same time, the team recognized the potential relevance of LLMs’ natural language capabilities for addressing their challenges related to emergency communication and social media dynamics. Several members expressed interest in generating realistic variations in agent behavior through persona-based prompting, viewing this as a way to capture community diversity and stress-test communication strategies. Notably, the team made no commitment to adapt simulation outputs into policy implementation, emphasizing that the collaboration was exploratory rather than prescriptive.